OpenStudy (anonymous):

-6 -(3-4)3 + 2(-2) _____________________ 3(-2 -1)

OpenStudy (anonymous):

\[-6-(-1)+(-4) \over 3(-3)\] = \[-6+1-4 \over -9\] = \[-9/-9 = 1\]

OpenStudy (anonymous):

-6-(-1)*3+2(-2) ---------------- 3(-3) = -6+3+-4 ---------- -9 = -3+-4 ------- -9 = -7 -- -9 =7/9

OpenStudy (anonymous):

^^ yeah I did a mistake, I didn't multiply by the three in the numerator!!